# fmeffects: An R Package for Forward Marginal Effects

### v1

@@ Conclusion
This paper introduces the R package fmeffects, the first software implementation of FMEs. We showcase the package functionality with an applied use case and discuss design choices and implications for future extensions. FMEs are a novel and versatile method for model-agnostic interpretations. They give us explanations in the form of: If we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$? FMEs provide comprehensible and actionable advice for stakeholders, including ones without machine learning expertise. We therefore hope that this package will work towards a more widespread adaption and usage of FMEs in practice.

### v2

@@ Conclusion
This paper introduces the R package fmeffects, the first software implementation of the theory surrounding FMEs. We showcase the package functionality with an applied use case and discuss design choices and implications for future extensions. FMEs are a versatile model-agnostic interpretation method and give us comprehensible model explanations in the form of: if we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$? FMEs equip stakeholders, including those without ML expertise, with the ability to understand feature effects for any model. We therefore hope that this package will work towards a more widespread adoption of FMEs in practice.
Software development is an ongoing process. As the theory surrounding FMEs evolves, so should the fmeffects package. As noted by Scholbeck et al. (2024), possible directions for future research include the development of techniques to better quantify extrapolation risk for the selection of step sizes; furthermore, the subgroup search for cAMEs is subject to uncertainties, which may be able to be quantified; and lastly, we may be able to spare computations by searching for representative FMEs, similar to prototype observations that are representative of clusters of observations (Tan et al., 2019). Future performance improvements may also be made via parallel computing, which at this point is only implemented for NLM computations.